Database last updated on Mon, 02 Feb 2026 23:10:54 +0000 by localhost
Instrument constant changes for the week before 2023299 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 20 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
aode
ds
dt
ed
fm
hg
rs
sl
uv
zs
299
SATURNA
#012
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.017)
50
50
1
2
36
15
1
6
52
23
(0 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B29923.012
1
#012
39
40